OpenStudy (anonymous):

How do I rewrite this function as a group of factors: f(x) = g^2 - 81 * I need the steps mainly

jimthompson5910 (jim_thompson5910):

hint: a^2 - b^2 = (a-b)(a+b) This is the difference of squares factoring rule

OpenStudy (anonymous):

I think the factors are (g + 9) (g- 9) but I do not know how to get to them.

jimthompson5910 (jim_thompson5910):

it might help to rewrite g^2 - 81 as g^2 - 9^2

jimthompson5910 (jim_thompson5910):

then use the rule: a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

I'll try to explain it the way my teacher taught me. So you know a polynomial function looks like this f(x)= x^2+x+n (where n is the numeral) in order to find the factors, you take n and multiply it with the coefficient of the x^2. Take for example, x^2+4x+4 4 x 1 = 4 (what, when added together gives you the middle coefficient of 4 and multiplies to become 4?) 2 and 2 so we have (x+2)(x+2) as factors. in this case, your function is f(g) = g^2 - 81 the middle coefficient is 0. you want to take 1 x -81 =-81 (what multiplied together gives you -81 but when added together gives you 0?) -9 & 9. So, the factors are: (g-9)(g+9) Does that make sense?

OpenStudy (anonymous):

So after I rewrite it as g^2 - 9^2 I just plug it into (a-b)(a+b) for my answer?

jimthompson5910 (jim_thompson5910):

correct, in this case, a = g and b = 9

jimthompson5910 (jim_thompson5910):

so.. a^2 - b^2 = (a-b)(a+b) turns into g^2 - 9^2 = (g-9)(g+9)